OpenStudy (anonymous):

multiply:

OpenStudy (anonymous):

\[\sqrt{3} * \sqrt{5}\]

Parth (parthkohli):

@rebeccaskell94 a practice for you

OpenStudy (anonymous):

uh oh beardo is here

OpenStudy (anonymous):

hahah thanks bro √15 I think

Parth (parthkohli):

Yes, correct

OpenStudy (anonymous):

yay thats what i thought too

OpenStudy (anonymous):

i got more ima post em :)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{a} \times \sqrt{b} = \sqrt{ab} }\)